OpenStudy (johnnydicamillo):

Use implicit differentiation to find the slope of the tangent line to the curve

OpenStudy (johnnydicamillo):

\[\frac{ y }{ x+4y} = x^8 - 9 \] at the point (1, -8/33)

OpenStudy (johnnydicamillo):

So I am going to use to the quotient rule on the first term

OpenStudy (johnnydicamillo):

\[\frac{ (x+4y)*y' - (y)(1+4y*y') }{ (x+4y)^2 }\]

OpenStudy (johnnydicamillo):

that is the first term, can I simplify that?

OpenStudy (freckles):

it looks good you can distribute and start gathering your y' terms on top

OpenStudy (johnnydicamillo):

@freckles will it look like this: \[\frac{ y'x+4 * yy'-y+4y^2 * yy' }{( x + 4y)^2}\]

OpenStudy (freckles):

close

OpenStudy (freckles):

\[\frac{ y'x+4 * yy'-y-4y^2 * y' }{( x + 4y)^2}\]

OpenStudy (johnnydicamillo):

okay, is that as far as I can simplify?

OpenStudy (freckles):

you can gather your y' terms if you want

OpenStudy (freckles):

you will need to do that to solve for y' later

OpenStudy (freckles):

have you differentiated the other side yet?

OpenStudy (johnnydicamillo):

so \[\frac{ y'x + 4*yy' - y -4y2 * y' }{ (x + 4y)^2 } = 8x^7 - 0\]

OpenStudy (freckles):

ok that is beautiful so far

OpenStudy (freckles):

so the only thing to do is put your y'

OpenStudy (freckles):

isolate your y'

OpenStudy (freckles):

you could undo that ugly division there on the left hand side by multiplying both sides by the (x+4y)^2

OpenStudy (freckles):

and then gather those y's on the left hand side and put everything that doesn't have a y' on the opposite side

OpenStudy (johnnydicamillo):

okay, let me give that a try

OpenStudy (johnnydicamillo):

\[y = x^3 + 4x^2y - 9x -36y\]

OpenStudy (johnnydicamillo):

now take the derivative of both sides?

OpenStudy (johnnydicamillo):

\[y' = 3x^2 + 8xy'-9-36y'\]

OpenStudy (freckles):

whoa whoa

OpenStudy (anonymous):

Wouldn't have used quotient rule. Would have used product rule.

OpenStudy (freckles):

you have already differentiated

OpenStudy (freckles):

something happened to all your y'

OpenStudy (freckles):

is this a different question you asking now

OpenStudy (freckles):

I'm confused

OpenStudy (johnnydicamillo):

I am too

OpenStudy (johnnydicamillo):

lol... I though you said to undo the division and multiply both sides

OpenStudy (freckles):

\[\frac{ y'x + 4*yy' - y -4y2 * y' }{ (x + 4y)^2 } = 8x^7 - 0 \\ y'x+4yy'-y-4y^2y'=8x^7(x+4y)^2 \] didn't you have this

OpenStudy (johnnydicamillo):

OH okay, I misunderstood, now let me see.

OpenStudy (freckles):

I multiply both sides by (x+4y)^2 to undo the division on the left

OpenStudy (freckles):

now you need to gather your terms on the left hand side that have y'

OpenStudy (johnnydicamillo):

I am having the trouble figuring the proper way to gather those terms, distribution?

OpenStudy (freckles):

factoring

OpenStudy (anonymous):

Since \( (x,y) =(1, -8/33)\)\[ \frac{ (x+4y)*y' - (y)(1+4y*y') }{ (x+4y)^2 }\\= \frac{ ((1)+4(-8/33))*y' - (-8/33)(1+4(-8/33)*y') }{ ((1)+4(-8/33))^2 } = (1)^8-9 \]

OpenStudy (freckles):

like you know ab+ac+e =a(b+c)+e

OpenStudy (johnnydicamillo):

okay, let me try

OpenStudy (freckles):

and I think that -9 isn't suppose to be there and that should be 8(1)^7 on the right

OpenStudy (johnnydicamillo):

\[y'(x + 4y)-y-4y^28y' = 8x^7\]

OpenStudy (johnnydicamillo):

is the first term correct?

OpenStudy (johnnydicamillo):

I forgot to append the rest on the second term

OpenStudy (freckles):

I think the 8 is a type-o \[y'(x+4y-4y^2)-y=8x^7(x+4y)^2\]

OpenStudy (freckles):

now you are solving for y' so add y on both sides then divide both sides by what y' is multiply multiplied by

OpenStudy (johnnydicamillo):

okay, let's see how messy this can get

OpenStudy (freckles):

it isn't too bad from here

OpenStudy (freckles):

just 2 steps to solve for y'

OpenStudy (johnnydicamillo):

\[y' = \frac{8x^7(x+4y^2)+y }{ x + 4y - 4y^2}\]

OpenStudy (johnnydicamillo):

I think the denominator can be factored

OpenStudy (johnnydicamillo):

or maybe not... will it cancel out?

OpenStudy (freckles):

hmmm not sure I can play it with it ...

OpenStudy (freckles):

and I think that square is missplaced on top

OpenStudy (freckles):

\[y'=\frac{8x^7(x+4y)^2+y}{x+4y-4y^2}\]

OpenStudy (freckles):

I can't really see a prettier way to write it

OpenStudy (freckles):

maybe you can write in a different form

OpenStudy (anonymous):

You do not need to find the implicit formula for \(y'\). You could have substituted in the point a long time ago to solve for the tangent line's slope.

OpenStudy (johnnydicamillo):

okay, assuming its correct, let's find the slope of the tangent line. So plug the x in corrent?

OpenStudy (johnnydicamillo):

@wio really?

OpenStudy (anonymous):

Yes, look what I wrote above.

OpenStudy (johnnydicamillo):

But that would be the slope is 8?

OpenStudy (anonymous):

No, you would need to solve for \(y'\).

OpenStudy (johnnydicamillo):

I am so lost.... I think I need to get rid of the Ys in the simplified equation

OpenStudy (jhannybean):

from what @wio had written (copy pasting) \[\frac{ (x+4y)*y' - (y)(1+4y*y') }{ (x+4y)^2 }\\= \frac{ ((1)+4(-8/33))*y' - (-8/33)(1+4(-8/33)*y') }{ ((1)+4(-8/33))^2 } = (1)^8-9\]All you'd have to d is factor out y' from the left hand side and shift all the other stuff over. It makes finding y' a lot easier .

OpenStudy (freckles):

\[y=(x^8-9)(x+4y) \\ y'=(8x^7-0)(x+4y)+(x^8-9)(1+4y') \\ y'=8x^7(x+4y)+x^8+4x^8y'-36y'-9 \\ y'-4x^8y'+36y'=8x^7(x+4y)+x^8-9 \\ y' (1-4x^8+36)=8x^7(x+4y)+x^8-9 \\ y'=\frac{8x^7(x+4y)+x^8-9}{ 37-4x^8}\] we could replace x^8-9 with y/(x+4y) \[y'=\frac{8x^7(x+4y)+\frac{y}{x+4y}}{1-4\frac{y}{x+4y}} = \frac{8x^7(x+4y)^2+y}{(x+4y)-4y} =\frac{8x^7(x+4y)^2+y}{x}\] So maybe we made a mistake somewhere I would have to go back through it -- you guys you are pluggin in the point without differentiating the right hand side

OpenStudy (anonymous):

\[ \frac{ y }{ x+4y} = x^8 - 9 \] at the point \((1, -8/33)\) I generally don't like the quotient rule, so I use the product rule.\[ \begin{split} &y = (x^8-9)(x+4y) \\ \implies &y'=(x^8-9)'(x+4y)+(x^8-9)(x+4y)'\\ \implies &y'=8x^7(x+4y)+(x^8-9)(1+4y') \end{split} \]It is at this point I would sub in \((x,y) = (1,-8/33)\):\[ y'=8(1)^7(1+4(-8/33))+((1)^8-9)(1+4y') \]Then it's a lot of simplification, and solving for \(y'\).

OpenStudy (johnnydicamillo):

but about the y' at the end of that?

OpenStudy (anonymous):

It's a variable. It will become easier to solve once you've gone through the order of operations.

OpenStudy (anonymous):

The purpose of implicit differentiation is to make it easier to solve for a derivative.

OpenStudy (freckles):

@johnnydicamillo I see the mistake now

OpenStudy (freckles):

it was in the quotient rule earlier

OpenStudy (freckles):

\[\frac{ (x+4y)*y' - (y)(1+4y') }{ (x+4y)^2 }\]

OpenStudy (freckles):

so after making that correction you would have \[y'=\frac{8x^7(x+4y)^2+y}{x+4y-4y} \]

OpenStudy (freckles):

since you already solved for y' you could also put your point into that

OpenStudy (freckles):

and yes the 4y-4y is 0

OpenStudy (freckles):

either way though

OpenStudy (johnnydicamillo):

okay to clear everything up \[y' = \frac{ (1)^7(1+4y)2+y }{ 1 }\]?

OpenStudy (johnnydicamillo):

scratch that. typos

OpenStudy (anonymous):

Suppose I even wanted to solve for \(y\) for whatever reason. I shouldn't do it in this problem for my own time, but if I wanted to do so then... \[ \frac{ y }{ x+4y} = x^8 - 9 \\ y = (x^8-9)(x+4y) =x(x^8-9)+4(x^8-9)y \\ y - 4(x^8-9)y =x(x^8-9)\\ y(1 - 4(x^8-9)) =x(x^8-9)\\ y = \frac{x(x^8-9)}{(1 - 4(x^8-9))} \]

OpenStudy (johnnydicamillo):

8/11?

OpenStudy (freckles):

\[y'=\frac{8(1)^7(1+4 \frac{-8}{33})^2+\frac{-8}{33}}{1}=8(1-\frac{32}{33})^2+\frac{-8}{33}=8(\frac{1}{33})^2-\frac{8}{33} =\frac{8}{33^2}-\frac{8}{33} \\ =\frac{8-8(33)}{33^2}\] I think this is right if i didn't make a mistake and yes I know this is kinda ugly I don't think that reduces to 8/11 looks negative to me also you can check my work there and finish it

OpenStudy (johnnydicamillo):

-256/1089

OpenStudy (johnnydicamillo):

yes!!!!!

OpenStudy (johnnydicamillo):

You are the best! Super messy problem